OpenStudy (anonymous):

Find the constant of variation for the relationship f(x)= 40x.

OpenStudy (anonymous):

im going to put your penis in my butt and then im going to poop on it and then when i get off off you u will pee out poop

OpenStudy (anonymous):

nasty

OpenStudy (anonymous):

umm.. im a girl soo ._.

OpenStudy (anonymous):

nasty people in the world

OpenStudy (the_fizicx99):

F(x) is another way of saying y, so you're given y = 40x, the formula for direct variation is y = kx where k, 40, is the constant making a table: x|y ----- 1 | 40 2| 80 3| 120 and so on, So if you divide both sides by x you have y/x= k Take a value from that table, 120/3 = 40 80/2 = 40 40/1 = 40

OpenStudy (the_fizicx99):

There's also indirect variation

OpenStudy (anonymous):

wow thanks much!.. n yeah there is..

OpenStudy (the_fizicx99):

yw

OpenStudy (anonymous):

sorry I did not know the answer!! :) im not that good